Angle A and angle B are supplementary. The measure of angle B is equal to four times the complement of angle A. What is the measure of angle A?
Answer by ewatrrr(24785)


Hi
A + B = 180
A + 4(90°-A) = 180° |B = four times the complement of angle A
- 3A = -180°
A = 60° AND B = 120°
CHECKING our answer***
60° + 4•30° = 180
Wish You the Best in your Studies.